Stock A has an annual expected return of 8%, a beta of .9, and a firm-specific volatility of 50% Stock B has an annual expected return of 9%, a beta of 1.3, and a firm-specific volatility of 40% The market has a standard deviation of 20%, and the risk-free rate is is 2%. Suppose we construct a portfolio built out of 50% stock A, 30% stock B, and 20% government t-bills.

What is the expected return of this portfolio? (in %, round to 1 decimal place) [b]

What is the beta of this portfolio? (round to 2 decimal places)
Answer #1

Stock A

Expected return on stock A = RA = 8%

Beta of stock A = βA = 0.9

Weight of stock A in the portfolio = wA = 50%

Stock B

Expected return on stock B = RB = 9%

Beta of stock B = βB = 1.3

Weight of stock B in the portfolio = wB = 30%

T-Bill

T-Bills are risk-free asset

Return on T-Bills = Risk free rate = RT = 2%

Beta of T-bills = βT = 0 [T-bills are risk-free]

Weight of T-bills in the portfolio = wT = 20%

Portfolio

Expected return on the portfolio is calculated using the formula:

E[RP] =wA*RA + wB*RB + wT*RT

Expected return on portfolio = E[RP] = (50%*8%) + (30%*9%) + (20%*2%) = 7.1%

Beta of the portfolio is calculated using the formula:

βP = wAA + wAB + wTT = (50%*0.9) + (30%*1.3) + (20%*0) = 0.84

Answers

a. Expected return on the portfolio = 7.1%

b. Beta of the portfolio = 0.84